OpenStudy (anonymous):

solve for n and simplify. -15/17=-25/-68 n

OpenStudy (saifoo.khan):

\[n = -\frac{12}{5}\]

OpenStudy (anonymous):

all the numbers in the fraction should be negative

OpenStudy (anonymous):

:)

OpenStudy (saifoo.khan):

\[\frac{-15}{17}=-\frac{-25}{-68} n\]

OpenStudy (anonymous):

17 is negative also

OpenStudy (saifoo.khan):

Oh, you wrote it incorrect then..

OpenStudy (saifoo.khan):

\[\frac{-15}{-17}=-\frac{-25}{-68} n\]

OpenStudy (saifoo.khan):

So, n = 12/5